Joint Cost
The Sun-Kissed Company manufactures two skin-care lotions, Soft Skin and Silken Skin, out of a joint process. The joint (common) costs incurred are $428,000 for a standard production run that generates 200,000 gallons of Soft Skin and 120,000 gallons of Silken Skin. Additional processing costs beyond the split-off point are $1.40 per gallon for Soft Skin and $0.90 per gallon for Silken Skin. Soft Skin sells for $2.40 per gallon while Silken Skin sells for $3.90 per gallon.

The Best Eastern Hotel chain has asked the Sun-Kissed Company to supply it with 240,000 gallons of Silken Skin at a price of $3.65 per gallon. Best Eastern plans to have the Silken Skin bottled in 1.5-ounce personal-use containers that are supplied in each of its hotel rooms as part of the complimentary personal products for guest use.

If Sun-Kissed accepts the order, it will save $0.05 per gallon in packaging of Silken Skin. There is sufficient excess capacity in Sun-Kissed's production system to handle just one more production run in order to have sufficient Silken Skin for this special order. However, the nature of the joint process always results in 200,000 gallons of Soft Skin and 120,000 gallons of Silken Skin. Also, the market for Soft Skin is saturated; hence, any additional sales of Soft Skin would take place at a price of $1.60 per gallon.

a. What is the profit normally earned on one production run of Soft Skin and Silken Skin?

Do not use negative signs with your answers.

Profit from one production run:

Soft skin Answer
Silken skin Answer
Total contribution margin Answer
Join production costs Answer
Total profit Answer

b. What is the incremental effect on overall income if the Sun-Kissed Company accepts the special order for Silken Skin?

Use a negative signs with your answer only to indicate a decrease in overall income. Otherwise do not use negative signs with your answers.

Additional contribution margin
Soft skin Answer
Silken skin Answer
Total Answer
Additional costs/lost profit:
Original Silken Skin contribution margin Answer
Joint production costs Answer
Increase (decrease) in overall income Answer
Answer #1

Answer page :01 Given data Sale potice soft skin PG = € 2.40 sale price Siliken askin = $ 3.90 per gallon Separate process coPage:02 profit from one production run. soft skin (A), 19.40-1.40) x 200,000 = 200,000 Siliken Skin (B) =( 3.90-0.90) x 12000page:03 Total = $ 80000 + $ 672000 = $ 752000 Additional cost / fost profit on additional skin contribution margin (D) ( 3.90Required page.!04 soft skin $ 200,000 Siliken skin $360000 Total contábution murgio 4560000 join production costs + 428000 To